Suppose a ball is drawn at random from a box containing three white and three black
balls. After a ball is drawn, it is then replaced and another drawn. What is the
probability that of the first four balls drawn, exactly two are white?
1
Expert's answer
2020-06-07T16:15:15-0400

"Probability\; of \;getting\; white\; ball=\frac{3}{6}\\
P(X=2)=C_{2}^{4}( \frac{3}{6} )^{2}(\frac{3}{6})^2=0.375"